OpenStudy (anonymous):

Domain of log equations questions. Help please! y=log(9-x^2) y=log(x^2 +4)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you can't take the log of a negative number you can't take the log of 0

jimthompson5910 (jim_thompson5910):

so for log(x), the domain is x > 0

jimthompson5910 (jim_thompson5910):

for log(9-x^2), the domain is 9-x^2 > 0 solve for x to get the answer for the first question

jimthompson5910 (jim_thompson5910):

for log(x^2 +4), the domain is x^2 +4 > 0 solve for x to get the answer for the second question

OpenStudy (anonymous):

so u get -3 or 3 is greater than x for the first one and get -2 or 2 for the second one right?

OpenStudy (anonymous):

we have to write the answer in interval notation so would I right (-3, infitity)?

jimthompson5910 (jim_thompson5910):

9-x^2 > 0 leads to -3 < x < 3 which in interval notation is (-3, 3)

jimthompson5910 (jim_thompson5910):

x^2 +4 is ALWAYS positive because x^2 is ALWAYS nonegative and you're adding on a positive number so x^2 +4 > 0 has the solution set of {x| x is any real number} So that in interval notation is (-infinity, infinity)

OpenStudy (anonymous):

I dont understand the first answer. How did it lead to -3 < x < 3?

jimthompson5910 (jim_thompson5910):

9-x^2 > 0 9 > x^2 x^2 < 9 sqrt(x^2) < sqrt(9) |x| < 3 -3 < x < 3

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

i just showed you

jimthompson5910 (jim_thompson5910):

in general sqrt(x^2) = |x| where x is any real number and |x| < k has the solution: -k < x < k where k is any positive number

OpenStudy (anonymous):

oh okay. culd u help with a few more?

jimthompson5910 (jim_thompson5910):

sure one more

OpenStudy (anonymous):

okay log(base 3)(x/x-1)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

finding the domain?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

and that's \(\Large \frac{x}{x-1}\) right?

OpenStudy (anonymous):

yes but there is a base 3 before it

jimthompson5910 (jim_thompson5910):

oh i'm just focusing on the stuff inside the log

jimthompson5910 (jim_thompson5910):

\(\Large \frac{x}{x-1}\) is only positive when... a) x and x-1 are both positive OR b) x and x-1 are both negative

jimthompson5910 (jim_thompson5910):

x and x-1 are both positive means x > 0 and x > 1 which compresses to x > 1

jimthompson5910 (jim_thompson5910):

x and x-1 are both negative means x < 0 and x < 1 which compresses to x < 0

jimthompson5910 (jim_thompson5910):

so... \(\Large \frac{x}{x-1}\) is only positive when... a) x > 1 OR b) x < 0

jimthompson5910 (jim_thompson5910):

how does that help us find the domain?

OpenStudy (anonymous):

not sure..

jimthompson5910 (jim_thompson5910):

well you can't take the log of a negative number or zero, so \(\Large \frac{x}{x-1}\) must be positive it's only positive when x > 1 or x < 0 as described above so the domain in interval notation is (-infinity, 0) U (1, infinity)